OpenStudy (anonymous):

Solve. 3x + 3 = 21 x = 5 x = 6* x = 18 x = 36 WILL GIVE MEDAL FOR CORRECT ANSWER!!!

OpenStudy (anonymous):

3x + 3 = 21 Subtract 3 from both sides: 3x = 18 Divide: x = 6

OpenStudy (otonogold):

The answer is 6. Plug in 6 3(6) + 3 =21 18+3=21